# Detecting Many-Body Scars from Fisher Zeros

### v2

@@ Abstract
Our work suggests that Fisher zeros carry more theoretical significance than previously recognized. The knowledge of Fisher zeros can aid the construction of parent Hamiltonians [57, 58, 59] to achieve novel quantum phases of matter. Non-unitary evolution in monitored quantum circuits [60, 61, 62] offers a natural way for realization and detection of complex $Z$. In addition, the rapidly improving capacity to measure the many-body states of Rydberg atoms [8, 9] promises probing Fisher zeros in finite systems. For example, one can extract the effective temperature $1/\beta_{r}$ by fitting the decay of the Loschmidt ratio for different states. Summing over these states gives the total $Z$ to reveal at which time it reaches zero. Given that Lee-Yang zeros have been detected by tuning external fields [63], it is reasonable to expect the experimental observation of Fisher zeros in the near future.
@@
Acknowledgments. We thank Xiwen Guan, Gaoyong Sun, Junsen Wang, and Li You for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12274126). E.Z. acknowledges the support from NSF Grant PHY-206419, and AFOSR Grant FA9550-23-1-0598. We conducted part of this work during the 2nd Conference on “Quantum Simulations of Fundamental Physics” and the International Workshop on “Quantum Systems with Novel Spatiotemporal Control” in the fall of 2024 in Shanghai.
Supplementary information for
@@
‘Detecting Many-Body Scars from Fisher Zeros’
I Fisher zeros of the $\bar{P}X\bar{P}$ model and the Fixed-point line
@@
to obtain $\beta_{i+1}$, and then use $\beta_{i+1}$ as the new starting point for iterative computation, continuing this process until $\beta$ reaches the fixed line $S_{0}=1$. Note that for the same $\beta_{i}$, all solutions $\beta_{i+1}$ satisfying Eq. S2 actually form a line. We choose the $\beta_{i+1}$ that is closest to $\beta_{i}$. In the main text, we presented the RG flow results for the matching between $L=32$ and $L=64$. To the left of the $S_{0}=1$ line, $S_{0}<1$, and $S_{0}=0$ is reached precisely at the Fisher zero. To the right of the $S_{0}=1$ line, $S_{0}>1$. The value of $|\ln S_{0}|$ reaches a minimum exactly at $S_{0}=1$ line, making this line the lowest point in the $|\ln S_{0}|$ landscape. Therefore, to more intuitively illustrate how the RG flow originates from either the Fisher zeros or large $\beta$ regions and approaches the $S_{0}=1$ line, we show in Fig. S3 the values of $|\ln S_{0}|$ for $L=64$, which clearly captures the “mountain range and water basin”-like structure discussed in the main text. In contrast to Ref. [2, 3], where the compare-$\beta$ RG flow terminates at an unique fix point ($\beta=0$), the newly defined RG flow here terminates on an extended fixed-point line . The flow toward Fisher zeros in Ref. [2, 3] exhibits ambiguous behavior for finite systems, whereas the newly defined RG flow always flows outward from the Fisher zeros, thereby avoiding this issue.
II Fisher zeros of $Z_{i}$ in the $\bar{P}X\bar{P}$ model
In the $\bar{P}X\bar{P}$ model, QMBS and SBETH not only exhibit different configurations of the Fisher zeros of $Z$ but also demonstrate distinct properties in the Fisher zeros of $Z_{i}$. At $g=1$, the zeros of all the $Z_{i}$ coincide with those of $Z$. As $g$ decreases and the blockade interaction is introduced, the zeros of most $Z_{i}$ remain close to the same position. This is also reflected in the dense lines of the solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$ in Fig. 2(a) of the main text. In contrast, in the QMBS region, $Z$ does not have Fisher zeros on the $\beta_{i}$-axis, implying that the zeros of $Z_{i}$ on the $\beta_{i}$ axis (if they exist) cannot coincide at the same position. Consequently, the dynamical behavior of each state will be completely different. We verify this conclusion by calculating the zeros of $Z_{i}$ for different states at $g=0$ with $L=32$. For a product state whose spin is configured with an up-up-down-down ($uudd$) pattern, we find that there are no zeros in a large region near the $\beta_{r}$ and $\beta_{i}$ axes [Fig. S4(a)]. We further consider different $T_{i}$ states (defined by a repeating pattern $uddd\cdots$, i.e. an up spin followed by $i-1$ down spins). For states closer to the FM state, we find that there are only one and two zeros exists near the axes in the $T_{32}$ and $T_{16}$ state, respectively [Fig. S4(b)(c)]. In contrast, for the scar state $T_{2}$ (or AF state), its zeros exhibit a clear tendency to approach the $\beta_{i}$ axis, and the positions differ significantly from those of the $T_{32}$ and $T_{16}$ states [Fig. S4(d)]. From this perspective, the differences in the zeros of $Z_{i}$ for all possible states in the Hilbert space are responsible for the zeros of the total $Z$ moving away from the $\beta_{i}$-axis.
@@ II Fisher zeros of $Z_{i}$ in the $\bar{P}X\bar{P}$ model
Figure S4: The Fisher zeros of boundary partition function $Z_{i}$ for different states in the $\bar{P}X\bar{P}$ model at $g=0$ (the $PXP$ model) with size $L=32$: (a) $uudd$ state, (b) $T_{32}$ state, (c) $T_{16}$ state, and (d) $T_{2}$ state (or AF state). The markings of lines and points are the same as in Fig. 2 of the main text. In (a), there are no zeros in the considered region, while (b)-(d) contain 1, 2, and 4 zeros, respectively. Their positions and the trend of approaching the $\beta_{i}$ axis differ significantly.
@@
III Fisher zeros of a lattice gauge field related model
@@
IV details on the Ising model with external fields
@@
We find that the off–$\beta_{i}$-axis Fisher zeros in the finite-size system still lie on the continuous line of zeros obtained analytically in the thermodynamic limit (Fig. S7), further supporting the validity of the results presented in the main text.

### v3

@@ Abstract
Our work suggests that Fisher zeros carry more theoretical significance than previously recognized. The knowledge of Fisher zeros can aid the construction of parent Hamiltonians [60, 61, 62] to achieve novel quantum phases of matter. Nonunitary evolution in monitored quantum circuits [63, 64, 65] offers a natural way for realization and detection of complex $Z$. In addition, the rapidly improving capacity to measure the many-body states of Rydberg atoms [8, 9] promises probing Fisher zeros in finite systems. For example, one can extract the effective temperature $1/\beta_{r}$ by fitting the decay of the Loschmidt ratio for different states. Summing over these states gives the total $Z$ to reveal at which time it reaches zero. Given that Lee-Yang zeros have been detected by tuning external fields [66], it is reasonable to expect the experimental observation of Fisher zeros in the near future.
@@
Acknowledgments—We thank Xiwen Guan, Gaoyong Sun, Junsen Wang, and Li You for helpful discussions. H.Z. is supported by the National Natural Science Foundation of China (Grant No. 12274126). E.Z. acknowledges the support from NSF Grant No. PHY-206419, and AFOSR Grant No. FA9550-23-1-0598. We conducted part of this work during the 2nd Conference on “Quantum Simulations of Fundamental Physics” and the International Workshop on “Quantum Systems with Novel Spatiotemporal Control” in Shanghai.
Data availability. The data that support the findings of this article are openly available [67].
I End Matter
Appendix A: Mountain range and water basin—To gain a deeper understanding of the location of Fisher zeros for QMBS and its qualitative difference from SBETH, we examine the analytical properties of $Z$ along the $\beta_{i}$ axis. Recall $Z=\sum_{i}Z_{i}$, where $\{|\psi_{i}\rangle\}$ form a complete basis and the zeros of $Z_{i}$ (DQPT) mark the singularities of $f_{i}=\ln|Z_{i}|$. In the case of SBETH [Fig. 4(a)], $f$ becomes nonanalytic over finite intervals (shaded regions), due to the collective effect of a broad set of states developing close-packed singularities in $f_{i}$. Interestingly, one of the singular points of the AFM state (red curve) is very close to the boundary of the nonanalytic region. In contrast, for the QMBS case both $f$ and $f^{\prime}$ remain smooth [Fig. 4(b)], different $Z_{i}$ exhibit different patterns of zeros in the complex $\beta$ plane [54], and the dynamics of $Z_{i}(\beta_{r}=0,\beta_{i})$ is also different for scar and nonscar states. In this case, the singularities of the non-scar FM state have little effect on $f$, while a singular point of the scar AFM state coincides with the inflection point of $f$ (shaded narrow region). The inflection point also corresponds to the projection of the lowest Fisher zero (see Fig. 2e) onto the $\beta_{i}$ axis. This “shadow effect” is analogous to how the rightmost Fisher zeros near the $\beta_{r}$ axis influence thermodynamic quantities such as specific heat [38].
Figure 4: Key differences between SBETH and QMBS in the analytical structure of $Z$ and the renormalization group (RG) flow. Left: $f=\ln|Z|$, $f^{\prime}\equiv df/d\beta_{i}$, and $f_{i}=\ln|Z_{i}|$ for the AFM (red) and FM (blue) states as functions of $\beta_{i}$ with $\beta_{r}=0$. In the SBETH case (a, $g=0.9$), $f$ becomes nonanalytic over a broad interval of $\beta_{i}$. In the QMBS case (b, $g=0$), both $f$ and $f^{\prime}$ evolve smoothly and $f$ exhibits only an inflection point. Right: the RG flows (arrows) for $g=0.9$ (c) and $g=0$ (d). The black lines connecting to the origin are the fixed points defined by $S_{0}=1$. The contour lines are for different values of $S_{0}$ at $L=64$.
The distinction between strong and weak ergodicity breaking can also be elucidated from renormalization group (RG). It has been noted that Fisher zeros form the boundary of RG flows on the complex $\beta$-plane [58, 59]. As shown in  [54], the “equipotential line” defined by $S_{0}=|Z(\beta_{r},t)/Z(0,0)|^{2}=1$ converges to a continuous line in $\beta$ plane as the system size $L\rightarrow\infty$, providing a convenient “high-temperature” (since it is connected to the origin $\beta=0$) fixed line for RG. The general trend of RG flows in Fig. 4 can be appreciated by a geological analogy: water flows toward the $S_{0}=1$ line (river in water basin) from near the Fisher zeros (mountain range). For SBETH, the Fisher zeros cross the $\beta_{i}$ axis, allowing the RG flow to reach the $S_{0}=1$ line throughout the $\beta$ plane [Fig. 4(c)]. The periodic return along the time ($\beta_{i}$) axis clearly violates the ETH. In the QMBS case, there exists a large plateau region between the zero line and the $\beta_{i}$ axis where the RG flows (not shown) are blocked by the “mountain range” of Fisher zeros and cannot reach the $S_{0}=1$ fixed line [Fig. 4(d)]. Thus the “mountain range” of Fisher zeros divides the $\beta$ plane into regions with distinctive long-time behaviors in $Z$, in agreement with the intuitive argument used earlier to motivate our conjecture.
Figure 5: Three-dimensional contour plots of $|\ln S_{0}|$ for $g=0.9$ (a) and $g=0$ (b).
We note that to the left of the $S_{0}=1$ line, $S_{0}<1$, and $S_{0}=0$ is reached precisely at the Fisher zero, while to the right of the $S_{0}=1$ line, $S_{0}>1$. The value of $|\ln S_{0}|$ reaches a minimum exactly at $S_{0}=1$ line, making this line the lowest point in the $|\ln S_{0}|$ landscape. Therefore, to more intuitively illustrate how the RG flow originates from either the Fisher zeros or large $\beta$ regions and approaches the $S_{0}=1$ line, we show in Fig. 5 the values of $|\ln S_{0}|$ for $L=64$, which clearly captures the “mountain range and water basinlike” structure as previously discussed.
@@ I End Matter
Supplementary information for ‘Detecting Many-Body Scars from Fisher Zeros’
II Fisher zeros of the $\bar{P}X\bar{P}$ model and the Fixed-point line
@@
to obtain $\beta_{i+1}$, and then use $\beta_{i+1}$ as the new starting point for iterative computation, continuing this process until $\beta$ reaches the fixed line $S_{0}=1$. Note that for the same $\beta_{i}$, all solutions $\beta_{i+1}$ satisfying Eq. S2 actually form a line. We choose the $\beta_{i+1}$ that is closest to $\beta_{i}$. In the main text, we presented the RG flow results for the matching between $L=32$ and $L=64$. In contrast to Ref. [58, 59], where the compare-$\beta$ RG flow terminates at an unique fix point ($\beta=0$), the newly defined RG flow here terminates on an extended fixed-point line . The flow toward Fisher zeros in Ref. [58, 59] exhibits ambiguous behavior for finite systems, whereas the newly defined RG flow always flows outward from the Fisher zeros, thereby avoiding this issue.
@@
Figure S3: The Fisher zeros of boundary partition function $Z_{i}$ for different states in the $\bar{P}X\bar{P}$ model at $g=0$ (the PXPPXP model) with size $L=32$: (a) uudduudd state, (b) $T_{32}$ state, (c) $T_{16}$ state, and (d) $T_{2}$ state (or AF state). The markings of lines and points are the same as in Fig. 2 of the main text. In (a), there are no zeros in the considered region, while (b)-(d) contain 1, 2, and 4 zeros, respectively. Their positions and the trend of approaching the $\beta_{i}$ axis differ significantly.
III Fisher zeros of $Z_{i}$ in the $\bar{P}X\bar{P}$ model
In the $\bar{P}X\bar{P}$ model, QMBS and SBETH not only exhibit different configurations of the Fisher zeros of $Z$ but also demonstrate distinct properties in the Fisher zeros of $Z_{i}$. At $g=1$, the zeros of all the $Z_{i}$ coincide with those of $Z$. As $g$ decreases and the blockade interaction is introduced, the zeros of most $Z_{i}$ remain close to the same position. This is also reflected in the dense lines of the solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$ in Fig. 2(a) of the main text. In contrast, in the QMBS region, $Z$ does not have Fisher zeros on the $\beta_{i}$-axis, implying that the zeros of $Z_{i}$ on the $\beta_{i}$ axis (if they exist) cannot coincide at the same position. Consequently, the dynamical behavior of each state will be completely different. We verify this conclusion by calculating the zeros of $Z_{i}$ for different states at $g=0$ with $L=32$. For a product state whose spin is configured with an up-up-down-down (uudduudd) pattern, we find that there are no zeros in a large region near the $\beta_{r}$ and $\beta_{i}$ axes [Fig. S3(a)]. We further consider different $T_{i}$ states (defined by a repeating pattern $uddd\cdots$, i.e. an up spin followed by i1i-1 down spins). For states closer to the FM state, we find that there are only one and two zeros exists near the axes in the $T_{32}$ and $T_{16}$ state, respectively [Fig. S3(b)(c)]. In contrast, for the scar state $T_{2}$ (or AF state), its zeros exhibit a clear tendency to approach the $\beta_{i}$ axis, and the positions differ significantly from those of the $T_{32}$ and $T_{16}$ states [Fig. S3(d)]. From this perspective, the differences in the zeros of $Z_{i}$ for all possible states in the Hilbert space are responsible for the zeros of the total $Z$ moving away from the $\beta_{i}$-axis.
@@ III Fisher zeros of $Z_{i}$ in the $\bar{P}X\bar{P}$ model
IV Fisher zeros of a lattice gauge field related model
@@
V details on the Ising model with external fields
@@
[…]